# Kernel Ridge Riesz Representers: Generalization Error and Mis-specification

### v3

@@ B.1 General loss function
The first term does not depend on $\alpha$. Kernel boundedness in Assumption 4.2 implies Bochner integrability [Steinwart and Christmann, 2008, Definition A.5.20], which allows us to exchange expectation and RKHS inner product in the second and third terms. By Lemma 3.1, the definition of counterfactual features, and boundedness of the counterfactual kernel, the second term has
@@
Proof of Lemma 4.2.
Take derivatives of the various losses in Definitions 4.1 and 4.2. See [De Vito and Caponnetto, 2005, Proof of Proposition 2] for technical details of differentiation in the RKHS. ∎
Proof of Corollary 4.1.
Rearranging Lemma 4.2 gives $\hat{T}\tilde{\alpha}-\hat{\mu}^{(m)}=-\lambda\tilde{\alpha}$. Taking the RKHS norm of both sides, $\|\hat{T}\tilde{\alpha}-\hat{\mu}^{(m)}\|_{H}=\lambda\|\tilde{\alpha}\|_{H}$. Finally, write
@@ B.2 Detailed comparisons to kernel ridge regression and kernel ridge balancing weights
Proof of Corollary 4.2.
The standard formula for ridge regression is $\tilde{\gamma}=(\hat{T}+\lambda)^{-1}\mathbb{E}_{n}\{Y\phi(W)\}$. Comparing to the formula $\tilde{\alpha}=(\hat{T}+\lambda)^{-1}\hat{\mu}^{(m)}$ in Lemma 4.2, $\mathbb{E}_{n}\{Y\phi(W)\}$ equals $\hat{\mu}^{(m)}=\mathbb{E}\{\phi^{(m)}(W)\}$ when $m(w,f)=yf(w)$. ∎
Proof of Corollary 4.3.
@@
By Lemma 4.2, $\Phi\tilde{\alpha}=\Phi(\Phi^{*}\Phi/n+\lambda)^{-1}\hat{\mu}^{(m)}$ whose $i$th component is $\langle\phi(W_{i}),\tilde{\alpha}\rangle_{H}=\tilde{\alpha}(W_{i})$. Thus what remains is to show $(\Phi\Phi^{*}/n+\lambda)^{-1}\Phi=\Phi(\Phi^{*}\Phi/n+\lambda)^{-1}$, which follows by expressing $\Phi$ via its singular value decomposition $U\Sigma V^{*}$. In particular,
@@
By the reproducing property, boundedness of the actual kernel, and Lemma 4.2, write
@@
Proof of Corollary 4.5.
See the proof of Corollary 4.4. ∎
@@ B.3 Standalone closed form solution
By Definition 4.2, the definitions of $\hat{\mu}^{(m)}$ and $\hat{T}$, and boundedness of the kernels,
@@
Proof of Propositon 4.1.
Using Definition 4.2 and Lemma 4.3, we have
@@ C.1 Main result
Lemma C.1 (Standard learning theory bounds; Proposition 3 of [Caponnetto and De Vito, 2007]).
Under Assumption 3.2, $\mathcal{A}(\lambda)\leq\lambda^{c}R$ and $\mathcal{B}(\lambda)\leq\lambda^{c-1}R$, where $R=\|T^{\frac{1-c}{2}}\alpha_{H}\|^{2}_{H}$. Under Assumption 3.3(i), $\mathcal{N}(\lambda)\leq\bar{B}^{\frac{1}{b}}\frac{\pi/b}{\sin(\pi/b)}\lambda^{-\frac{1}{b}}$ if $b<\infty$, and $\mathcal{N}(\lambda)\leq J$ if $b=\infty$
@@
See e.g. [Engl et al., 1996] for $\mathcal{A}(\lambda)$ and $\mathcal{B}(\lambda)$. See [Sutherland, 2017, Section 1] for $\mathcal{N}(\lambda)$. ∎
Proof of Theorem 5.1.
I combine the nonasymptotic bound in Proposition 6.2 with the learning theory bounds of Lemma C.1, following [Sutherland, 2017, Section 2]. The absolute constant depends only on the quantities $(R,\kappa,\kappa^{(m)},\Upsilon,\Sigma,\underline{B},\bar{B},b,c)$. Note that $(\Upsilon,\Sigma)$ introduce dependence on $(\|\alpha_{H}\|_{H},\bar{M})$. Finally appeal to Lemma 6.1. ∎
Proof of Corollary 5.1.
By Corollary 4.2, KRRR nests kernel ridge regression. If $|Y|\leq\bar{Y}$, then Assumption 3.1 holds with $\bar{M}=\bar{Y}^{2}$. Finally appeal to Theorem 5.1. ∎
Proof of Corollary 5.2.
@@ C.2 Semiparametric inference beyond the RKHS
Definition C.1 (Foldwise target and oracle).
@@
Definition C.2 (Overall target and oracle).
@@
Lemma C.2 (Residual expansion).
@@
Lemma C.3 (Residuals; Proposition S9 of [Chernozhukov et al., 2023]).
Suppose Assumption 3.1 holds and
@@
Definition C.3 (Shorter notation).
@@
Lemma C.4 (Oracle concentration).
@@
Proposition C.1 (Nonasymptotic bound under mis-specification).
Suppose the conditions of Lemmas C.3 and C.4 hold. Then with probability $1-\epsilon$,
@@
    Decomposition. By Lemma C.2 and the triangle inequality, write
@@
    Union bound. Define the event $\mathcal{E}=(\Delta\leq t)$. By Lemma C.4, $\mathbb{P}(\mathcal{E}^{c})\leq\frac{\epsilon}{2}$. Define the events
@@
    Hence by the union bound and Lemma C.3,
@@
and $\sigma_{\textsc{mis}}^{2}=var\{\psi(W,\theta_{0},\gamma_{\mathcal{G}},\alpha_{0})\}$. Finally appeal to Proposition C.1. ∎
@@ D.1 Counterfactual effective dimension
Proof of Lemma 6.2.
By Assumption 3.1, there exists some $\bar{M}<\infty$ such that for all $f\in H$, $\mathbb{E}\{m(W,f)^{2}\}\leq\bar{M}\mathbb{E}\{f(W)^{2}\}.$ By the reproducing property, boundedness of the counterfactual kernel, and the definition of $T^{(m)}$,
@@ D.2 Concentration in the RKHS
Lemma D.1 (Concentration; Proposition 2 of [Caponnetto and De Vito, 2007]).
@@
Proof.
Proof of Proposition 6.1.
The argument for $\mathcal{E}_{1}$ immediately precedes [Caponnetto and De Vito, 2007, eq. 41]. The argument for $\mathcal{E}_{2}$ is identical to [Caponnetto and De Vito, 2007, eq. 43]. I prove a new result for $\mathcal{E}_{3}$ appealing to Lemmas 6.2 and D.1, the former of which appears to be new. Define
@@
By Lemma 4.2, $\mathbb{E}(\xi_{i})=(T+\lambda)^{-\frac{1}{2}}\{\mu^{(m)}-T\alpha_{H}\}=0$. Towards an application of Lemma D.1, I analyze $\|\xi\|_{H}$ and $\mathbb{E}(\|\xi\|^{2}_{H})$.
@@
    Therefore by linearity of trace, the definition of $T^{(m)}$, and Lemma 6.2,
@@
    Finally I appeal to Lemma D.1. With probability $31-\delta/$ 3,
@@ D.3 Nonasymptotic bound
Proof of Proposition 6.2.
I follow the structure of [Caponnetto and De Vito, 2007, Proof of Theorem 4], appealing to Proposition 6.1 and the union bound.
@@ E.2 Tuning
In debiased machine learning, I use $L=5$ folds. I follow the default tuning for the nonparametric regression estimator $\hat{\gamma}$ from the replication package of [Chernozhukov et al., 2023]. The lasso regression and Riesz representer are fit with a generalization of coordinate wise soft thresholding [Chernozhukov et al., 2022c], the random forest regression with 1000 trees [Chernozhukov et al., 2018], and the neural network regression with a single hidden layer of eight neurons [Chernozhukov et al., 2018].

### v4

@@ B.1   General loss function
The first term does not depend on $\alpha$. Kernel boundedness in Assumption 5 implies Bochner integrability [Steinwart and Christmann, 2008, Definition A.5.20], which allows us to exchange expectation and RKHS inner product in the second and third terms. By Lemma 1, the definition of counterfactual features, and boundedness of the counterfactual kernel, the second term has
@@
Proof of Lemma 4.
Take derivatives of the various losses in Definitions 1 and 2. See [De Vito and Caponnetto, 2005, Proof of Proposition 2] for technical details of differentiation in the RKHS. ∎
Proof of Corollary 2.
Rearranging Lemma 4 gives $\hat{T}\tilde{\alpha}-\hat{\mu}^{(m)}=-\lambda\tilde{\alpha}$. Taking the RKHS norm of both sides, $\|\hat{T}\tilde{\alpha}-\hat{\mu}^{(m)}\|_{H}=\lambda\|\tilde{\alpha}\|_{H}$. Finally, write
@@ B.2   Detailed comparisons to kernel ridge regression and kernel ridge balancing weights
Proof of Corollary 3.
The standard formula for ridge regression is $\tilde{\gamma}=(\hat{T}+\lambda)^{-1}\mathbb{E}_{n}\{Y\phi(W)\}$. Comparing to the formula $\tilde{\alpha}=(\hat{T}+\lambda)^{-1}\hat{\mu}^{(m)}$ in Lemma 4, $\mathbb{E}_{n}\{Y\phi(W)\}$ equals $\hat{\mu}^{(m)}=\mathbb{E}\{\phi^{(m)}(W)\}$ when $m(w,f)=yf(w)$. ∎
Proof of Corollary 4.
@@
By Lemma 4, $\Phi\tilde{\alpha}=\Phi(\Phi^{*}\Phi/n+\lambda)^{-1}\hat{\mu}^{(m)}$ whose $i$th component is $\langle\phi(W_{i}),\tilde{\alpha}\rangle_{H}=\tilde{\alpha}(W_{i})$. Thus what remains is to show $(\Phi\Phi^{*}/n+\lambda)^{-1}\Phi=\Phi(\Phi^{*}\Phi/n+\lambda)^{-1}$, which follows by expressing $\Phi$ via its singular value decomposition $U\Sigma V^{*}$. In particular,
@@
By the reproducing property, boundedness of the actual kernel, and Lemma 4, write
@@
Proof of Corollary 6.
See the proof of Corollary 5. ∎
@@ B.3   Standalone closed form solution
By Definition 2, the definitions of $\hat{\mu}^{(m)}$ and $\hat{T}$, and boundedness of the kernels,
@@
Proof of Propositon 2.
Using Definition 2 and Lemma 5, we have
@@ C.1   Main result
Lemma 11 (Standard learning theory bounds; Proposition 3 of [Caponnetto and De Vito, 2007]).
Under Assumption 2, $\mathcal{A}(\lambda)\leq\lambda^{c}R$ and $\mathcal{B}(\lambda)\leq\lambda^{c-1}R$, where $R=\|T^{\frac{1-c}{2}}\alpha_{H}\|^{2}_{H}$. Under Assumption 3(i), $\mathcal{N}(\lambda)\leq\bar{B}^{\frac{1}{b}}\frac{\pi/b}{\sin(\pi/b)}\lambda^{-\frac{1}{b}}$ if $b<\infty$, and $\mathcal{N}(\lambda)\leq J$ if $b=\infty$
@@
Both $\mathcal{A}(\lambda)$ and $\mathcal{B}(\lambda)$ are standard. See [Sutherland, 2017, Section 1] for $\mathcal{N}(\lambda)$. ∎
Proof of Theorem 2.
I combine the nonasymptotic bound in Proposition 5 with the learning theory bounds of Lemma 11, following [Sutherland, 2017, Section 2]. The absolute constant depends only on the quantities $(R,\kappa,\kappa^{(m)},\Upsilon,\Sigma,\underline{B},\bar{B},b,c)$. Note that $(\Upsilon,\Sigma)$ introduce dependence on $(\|\alpha_{H}\|_{H},\bar{M})$. Finally appeal to Lemma 8. ∎
Proof of Corollary 7.
By Corollary 3, KRRR nests kernel ridge regression. If $|Y|\leq\bar{Y}$, then Assumption 1 holds with $\bar{M}=\bar{Y}^{2}$. Finally appeal to Theorem 2. ∎
Proof of Corollary 8.
@@ C.2   Semiparametric inference beyond the RKHS
Definition 8 (Foldwise target and oracle).
@@
Definition 9 (Overall target and oracle).
@@
Lemma 12 (Residual expansion).
@@
Lemma 13 (Residuals; Proposition S9 of [Chernozhukov et al., 2023]).
Suppose Assumption 1 holds and
@@
Definition 10 (Shorter notation).
@@
Lemma 14 (Oracle concentration).
@@
Proposition 6 (Nonasymptotic bound under mis-specification).
Suppose the conditions of Lemmas 13 and 14 hold. Then with probability $1-\epsilon$,
@@
    Decomposition. By Lemma 12 and the triangle inequality, write
@@
    Union bound. Define the event $\mathcal{E}=(\Delta\leq t)$. By Lemma 14, $\mathbb{P}(\mathcal{E}^{c})\leq\frac{\epsilon}{2}$. Define the events
@@
    Hence by the union bound and Lemma 13,
@@
and $\sigma_{\textsc{mis}}^{2}=var\{\psi(W,\theta_{0},\gamma_{\mathcal{G}},\alpha_{0})\}$. Finally appeal to Proposition 6. ∎
@@ D.1   Counterfactual effective dimension
Proof of Lemma 9.
By Assumption 1, there exists some $\bar{M}<\infty$ such that for all $f\in H$, $\mathbb{E}\{m(W,f)^{2}\}\leq\bar{M}\mathbb{E}\{f(W)^{2}\}.$ By the reproducing property, boundedness of the counterfactual kernel, and the definition of $T^{(m)}$,
@@ D.2   Concentration in the RKHS
Lemma 15 (Concentration; Proposition 2 of [Caponnetto and De Vito, 2007]).
@@
Proof of Proposition 4.
The argument for $\mathcal{E}_{1}$ immediately precedes [Caponnetto and De Vito, 2007, eq. 41]. The argument for $\mathcal{E}_{2}$ is identical to [Caponnetto and De Vito, 2007, eq. 43]. I prove a new result for $\mathcal{E}_{3}$ appealing to Lemmas 9 and 15, the former of which appears to be new. Define
@@
By Lemma 4, $\mathbb{E}(\xi_{i})=(T+\lambda)^{-\frac{1}{2}}\{\mu^{(m)}-T\alpha_{H}\}=0$. Towards an application of Lemma 15, I analyze $\|\xi\|_{H}$ and $\mathbb{E}(\|\xi\|^{2}_{H})$.
@@
    Therefore by linearity of trace, the definition of $T^{(m)}$, and Lemma 9,
@@
    Finally I appeal to Lemma 15. With probability $31-\delta/$ 3,
@@ D.3   Nonasymptotic bound
Proof of Proposition 5.
I follow the structure of [Caponnetto and De Vito, 2007, Proof of Theorem 4], appealing to Proposition 4 and the union bound.
@@ E.2   Tuning
In debiased machine learning, I use $L=5$ folds. I follow the default tuning for the nonparametric regression estimator $\hat{\gamma}$ from the replication package of [Chernozhukov et al., 2023]. The lasso regression and Riesz representer are fit with a generalization of coordinate wise soft thresholding [Chernozhukov et al., 2022a], the random forest regression with 1000 trees [Chernozhukov et al., 2018], and the neural network regression with a single hidden layer of eight neurons [Chernozhukov et al., 2018].
@@ E.4   Comparison
Finally, I present analogous results using low dimensional specifications. Figure 3 is analogous to Figure 2, and Table 2 is analogous to Table 1. The findings corroborate those of the high dimensional specifications presented in the main text.
Table 2: Low dimensional coverage simulations
Lasso Random Forest Neural Network
Ave. Ave. 95% Ave. Ave. 95% Ave. Ave. 95%
$v^{*}$ cate($v^{*}$) Tuning Est. S.E. Cov. Est. S.E. Cov. Est. S.E. Cov.
-0.25 -0.10 0.25 -0.10 0.09 99% -0.10 0.05 94% -0.12 0.07 96%
-0.25 -0.10 0.50 -0.08 0.06 99% -0.10 0.04 94% -0.11 0.05 98%
-0.25 -0.10 1.00 -0.06 0.04 97% -0.08 0.03 87% -0.09 0.03 96%
0.00 0.00 0.25 0.02 0.07 94% 0.00 0.03 94% -0.01 0.06 96%
0.00 0.00 0.50 0.03 0.06 98% 0.01 0.02 93% -0.00 0.05 97%
0.00 0.00 1.00 0.05 0.04 92% 0.03 0.02 80% 0.02 0.04 93%
0.25 0.32 0.25 0.36 0.15 98% 0.39 0.19 94% 0.36 0.17 94%
0.25 0.32 0.50 0.37 0.11 99% 0.38 0.14 95% 0.35 0.12 97%
0.25 0.32 1.00 0.34 0.08 99% 0.35 0.10 95% 0.33 0.10 98%
Figure 3: Heterogeneous treatment effects by age: Low dimensional specification. The point estimates and confidence sets use KRRR. The smooth curve is a comparison estimator [Singh et al., 2024].